If Stanford accepts 3 transfer students per year out of 30,000 applicants in California what is the probability that out of 15,400 applicants from southern California only one or zero get in?
Answer by Boreal(15235)

Poisson with proportional to number, random and theoretically infinite or very, very large. Probability is 1 in 10,000 or 0.0001
np=lambda=15400*0.0001=1.54= lambda. When n goes to infinity, p goes to 0, but np is a constant NE to 0, then use this.
e^(-lambda)^lambda^x/x!
e^(-1.54)1.54^1/=0.3301 for 1
e^(-1.54)=0.2144 for 0
so 0 or 1 has a probability of 0.5445.